OpenStudy (anonymous):

Solve 3x2 - 12x + 2 = 0 by using a method different from the one you used in Part B. Show the steps of your work can you solve using quadratic formula please!! @TuringTest @dan815 @nincompoop

OpenStudy (anonymous):

@amistre64 @ganeshie8

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

referring to a generic quadratic equation: a*x^2+b*x+c=0, the roots, are given by the subsequent formula (quadratic formula): \[\Large x = \frac{{ - b \pm \sqrt {{b^2} - 4ac} }}{{2a}}\]

OpenStudy (anonymous):

okay.

OpenStudy (michele_laino):

now in our case , we have: a=3, b=-12, c=2

OpenStudy (anonymous):

okayy

OpenStudy (michele_laino):

and the roots are: \[\Large \begin{gathered} x = \frac{{ - b \pm \sqrt {{b^2} - 4ac} }}{{2a}} = \hfill \\ \hfill \\ = \frac{{12 \pm \sqrt {{{12}^2} - 4 \times 3 \times 2} }}{{2 \times 3}} = ... = \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

so i solve that

OpenStudy (anonymous):

i got 16?

OpenStudy (anonymous):

12+-square root 144-24/6

OpenStudy (michele_laino):

I got this: \[\Large \begin{gathered} \frac{{12 \pm \sqrt {{{12}^2} - 4 \times 3 \times 2} }}{{2 \times 3}} = \frac{{12 \pm \sqrt {144 - 24} }}{6} = \hfill \\ \hfill \\ = \frac{{12 \pm \sqrt {120} }}{6} = \frac{{12 \pm 2\sqrt {30} }}{6} = \frac{{6 \pm \sqrt {30} }}{3} \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

holy thats confusing

OpenStudy (michele_laino):

we can write this: \[\sqrt {120} = \sqrt {4 \times 30} = \sqrt {{2^2} \times 30} = 2\sqrt {30} \]

OpenStudy (anonymous):

so what would be my answer lol im confused xD?

OpenStudy (michele_laino):

more explanation: \[\frac{{12 \pm \sqrt {120} }}{6} = \frac{{12 \pm 2\sqrt {30} }}{6} = \frac{{2\left( {6 \pm \sqrt {30} } \right)}}{6} = \frac{{6 \pm \sqrt {30} }}{3}\] the roots of your equation are: \[{x_1} = \frac{{6 + \sqrt {30} }}{3},\quad {x_2} = \frac{{6 - \sqrt {30} }}{3}\]

OpenStudy (anonymous):

thats my answer correcT?

OpenStudy (michele_laino):

yes! nevertheless you have to simplify the expression that you have got

OpenStudy (anonymous):

Okay! can you help me on one more question?

OpenStudy (michele_laino):

yes! please wait, someone is calling me to my phone